OpenStudy (anonymous):

Help i got stuck >:| ((2x-5/x-3)/(x-3))-2 = 3/x+3 I moved 2x-5/x-3 to the other side. I think the common denominator would be (x-3)(x+3).....

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{\frac{2x-5}{x-3}}{x-3}=\frac{3}{x+3}}\) like this ?

OpenStudy (anonymous):

Help i got stuck >:| ((2x-5)/(x-3))-2 = 3/x+3 Sorry about that it should be like this